OpenStudy (rainbow_rocks03):

Find the discount. regular price = $349; discount rate = 1/10 A. $3.49 B. $34.90 C. $314.10 D. $3490

OpenStudy (rainbow_rocks03):

@campbell_st

OpenStudy (australopithecus):

Multiply 349 by (1/10) you can use the rule: \[a*(\frac{c}{d}) = \frac{a*c}{d}\]

OpenStudy (rainbow_rocks03):

so my answer is B

OpenStudy (rainbow_rocks03):

thx

OpenStudy (australopithecus):

Wait

OpenStudy (rainbow_rocks03):

oh

OpenStudy (australopithecus):

Do they define what they mean by discount rate?

OpenStudy (rainbow_rocks03):

huh?

OpenStudy (australopithecus):

Well 1/10 is 10% so when you multiply that by 349 you get 10% of 349 so you pay 10% of the original price which seems kind of crazy Another way of interpreting it is that you get 10% off of 349 in that case you would have to subtract 1 by 1/10 then multiply by 349 1 - 1/10 gives you 90% so after multiplying you would get 10% off of 349 which seems like a more reasonable discount.

OpenStudy (rainbow_rocks03):

So my answer is B?

OpenStudy (australopithecus):

I would assume that it is 10% off the original price because paying 10% of the original price is just bonkers

OpenStudy (australopithecus):

do you follow?

OpenStudy (australopithecus):

are you still here?